# SOLD: Reinforcement Learning with Slot Object-Centric Latent Dynamics

### v1

@@ 4 Results
Figure 3: Open-loop predictions on the tasks PickAndPlace-Distinct (top) and Hammer (bottom). Starting from a single context frame, our model predicts the next 50 frames by propagating the individual slot representations forward without access to any intermediate images.
We evaluate SOLD on a suite of robotic manipulation tasks, designed to test for both complex relational reasoning as well as low-level manipulation capabilities. Further, we apply our method to environments that are not designed as object-centric tasks. Specifically, we test on two environments from Meta-World (Yu et al., 2019) and DM-Control (Tassa et al., 2018), respectively.
@@
We design the experiments to achieve two main objectives: first, to specifically assess the impact of the object-centric paradigm in our method by comparing it to a baseline that replaces the object-centric encoder-decoder modules with a standard convolutional architecture (Ours w/o OCE); and second, to evaluate our approach against the best available competitor from the literature by benchmarking it against DreamerV3 (Hafner et al., 2023), a state-of-the-art MBRL algorithm known for its strong performance across a wide range of tasks. Here, we choose the 12 million parameter version, to match the parameter count of our own model. Further details about the baselines are provided in Appendix C.
@@
We introduce a suite of eight object-centric robotic control environments designed to test both relational reasoning and manipulation capabilities. These environments feature two types of problems: Reach tasks, where the agent must identify a target and move the end-effector to its location, and manipulation tasks (Push and PickAndPlace), where the agent identifies a target block and moves it to a designated goal. The action-space is 4-dimensional, where the first three components represent the desired movement direction of the end effector, and the fourth controls the gripper. On the Reach and Push tasks, commands to the gripper are ignored, with the gripper fixed in a closed configuration, as gripping is not required to solve these tasks. To test varying levels of relational reasoning difficulty, we design the following tasks:
@@
For the Reach task, we also introduce two more challenging variants:
  • Specific-Relative    Instead of reaching for the red object, the goal is to reach the reddest object, determined by the perceptual CIEDE2000 (Sharma et al., 2005) distance.
  • Distinct-Groups    The environment always contains 5 targets, and the goal is to reach the one that appears only once.
On these two additional reach tasks, we reuse the SAVi models that were pre-trained for Reach-Specific and Reach-Distinct, respectively without modification. Further details about these environments are provided in Appendix D.
In addition to these tasks, we evaluate our approach on environments not originally designed for object-centric learning to investigate its generalizability. We include the Button-Press and Hammer tasks from the Meta-World benchmark (Yu et al., 2019), both featuring objects with complex shapes and textures, testing the model’s ability to handle more diverse visual inputs. Finally, we assess our method on the Cartpole-Balance and Finger-Spin environments from the DM-Control suite (Tassa et al., 2018), which represent vastly different domains not typically associated with object-centric approaches. An overview of all studied environments is given in Figure 1.
@@ 4.1 Object-centric Dynamics Learning
The object-centric representations learned by SAVi can be seen in the context-frames of Figure 3. The slots effectively decompose the visual scene, with most slots representing distinct objects and three slots capturing different parts of the respective robots. This part-whole segmentation demonstrates that the slots can meaningfully identify separate parts of a larger object, representing the gripper jaws in the first example and different parts of the kinematic chain of the Sawyer robot in the second. Notably, the sharp mask predictions show that each slot isolates information about the specific object it represents (see also Section E). This property is crucial for object-centric behavior learning, as it enables subsequent components to reason about task-relevant objects while ignoring irrelevant information.
Further, the open-loop predictions visualized in Figure 3 demonstrate that the object decomposition learned by the SAVi model is preserved throughout the prediction sequence. On the PickAndPlace-Distinct task, the movement of the robot and the blocks is predicted with high accuracy, showcasing the model’s ability to capture complex physical interactions. Furthermore, the model effectively handles occlusions, as evidenced by the continued precise prediction of the spherical red target. In the second example, the slots capture the intricate shape of the hammer and nail-box. The predictions remain reliable over a long horizon, even during interactions between the robot and the hammer, and between the hammer and nail.
Figure 4: Final success rates across the eight evaluated environments. Overall, SOLD compares favorably to DreamerV3, showing slight improvements on the Specific tasks and a significant performance boost on the Distinct variants.
@@ 4.2 Behavior Learning
To assess SOLD’s performance across our suite of robotic control tasks, we train each method three times with different random seeds on each environment. The final success rate achieved by each method is shown in Figure 4. SOLD consistently outperforms the non-object-centric baseline, often by a significant margin, highlighting the effectiveness of object-centric representations for these problems. While the non-object-centric baseline is able to make progress on the Specific variants of the tasks, it struggles to perform the relational reasoning required to solve the Distinct versions. As a result, we decided excluded it from our evaluation of the two more advanced versions of the Reach tasks. When compared to the state-of-the-art method DreamerV3, our model shows competitive or superior performance. This advantage is particularly pronounced for tasks that require relational reasoning between objects, namely the Distinct variants and the Specific-Relative task.
These results confirm our hypothesis that learning a structured latent representation in the world model benefits downstream tasks that require object reasoning, which are common in robotics contexts. Beyond that, when examining the performance over the course of training, as shown in Figure 5, we observe additional advantages. Even after accounting for the samples used during pre-training, our method consistently outperforms the highly sample-efficient DreamerV3 baseline, learning quickly and achieving high returns with minimal experience. This enhanced sample efficiency is observed across all tasks, highlighting the utility of the structured latent space to behavior models.
Figure 5: Achieved return over the training duration on our eight benchmark environments. The dotted vertical line indicates the offset of our method to account for the data used during pre-training.
@@
In Figure 6, we visualize an extract of the slot history via the image reconstructions visualized in the first row. To visualize the attention pattern of the actor in the current (rightmost) time-step we multiply the attention scores with the masks of the respective objects and show them overlaid with the RGB reconstructions and as an individual colormap in the second and third row respectively. This visualization shows the Push-Specific task and we find that the model discovers task relevant objects automatically, ignoring information stored in slots that represent distractor objects across all time-steps, while the robot and green cube receive the most attention. While the recency bias induced by ALiBi is clearly visible, we find that the model learns to pierce through it when necessary, attending to the target, which has been occluded for 15 time-steps in the last time-step where it was visible. We see that the model is able to focus on task-relevant information effectively, even when reasoning over long time sequences is required.
@@
One common limitation of prior works is that object-centric models are pretrained on sequences with random behaviors but then frozen during training. This restricts their applicability to tasks where the state distributions encountered by random and successful policies are similar. With the PickAndPlace tasks, we explicitly violate this assumption since random behaviors are highly unlikely to pick and lift blocks off the table. Hence, SAVi has not seen blocks in the air, a configuration that will necessarily be reached by performant policies. Figure 7 illustrates the need to continually fine-tune the object-centric encoder-decoder model. While the fine-tuned SAVi model is able to reconstruct the lifted block accurately, the frozen variant fails to decode this configuration correctly. The target block effectively dissolves when lifted, hindering the discovery of behaviors associated with such states.
Generalization to Non-Object-Centric Environments
While it is commonplace to evaluate object-centric methods on environments and datasets that naturally lend themselves to such decompositions, we aim to showcase the potential of our methods to generalize beyond this setting. To this end, we tested our method on Meta-World to assess its performance on tasks involving objects with complex shapes and colors, rather than simple uni-colored objects. Our method converges to success rate of $100\%$ on both the Button-Press and Hammer task. Further, we test on DM-Control to evaluate generalization to problems that are vastly different and where reasoning over objects and interactions is less pronounced. On the Cartpole-Balance and Finger-Spin tasks, we reach returns of $497$ and $645$, respectively. We provide details about the environment decomposition and dynamics prediction for all four tasks in Section E in the Appendix.
@@ 5 Related Work
In recent years, the field of unsupervised object-centric representation learning from images and videos has gained significant attention (Yuan et al., 2023). Most existing methods follow an encoder-decoder framework with a structured bottleneck composed of $N$ latent vectors called slots, where each of these slots binds to a different object in the input image. Slot-based methods have been widely applied for images (Burgess et al., 2019; Engelcke et al., 2020; Locatello et al., 2020; Singh et al., 2021; Engelcke et al., 2021; Singh et al., 2023; Biza et al., 2023) and videos (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022; Bao et al., 2022; Zoran et al., 2021; Kabra et al., 2021; Creswell et al., 2021). However, despite their impressive performance on synthetic datasets, they often fail to generalize to visually complex scenes. To overcome this limitation, recent methods propose introducing some weak supervision (Elsayed et al., 2022; Bao et al., 2023), levering large self-supervised pretrained encoders  (Seitzer et al., 2023; Zadaianchuk et al., 2024; Aydemir et al., 2023; Kakogeorgiou et al., 2024), or using diffusion models as slot decoders (Jiang et al., 2023; Wu et al., 2023b; Singh et al., 2024).
@@
Object-centric video prediction aims to understand the object dynamics in a video sequence with the goal of anticipating how these objects will move and interact with each other in future time steps. With this end, multiple methods propose to model and forecast the object dynamics using different architectures, including RNNs (Zoran et al., 2021; Nakano et al., 2023) transformers (Wu et al., 2023a; Villar-Corrales et al., 2023; Song et al., 2023; Gandhi et al., 2024; Daniel & Tamar, 2024; Nguyen et al., 2024; Petri et al., 2024) or state-space models (Jiang et al., 2024), achieving an impressive prediction performance on synthetic video datasets and learning representations that can help solve downstream tasks that require reasoning about objects properties and relationships (Wu et al., 2023a; Petri et al., 2024).
Model-based Reinforcement Learning
Model-based reinforcement learning approaches aim to improve sample efficiency by learning environment dynamics. PlaNet (Hafner et al., 2019) introduced a latent dynamics model for efficient planning, while the Dreamer family (Hafner et al., 2020; 2021; 2023) incorporated this into an actor-critic framework. DreamerV2 and DreamerV3 introduced further improvements like categorical latent states and robustness techniques. DreamerV3 has shown superior performance in visual control tasks compared to model-free approaches, but uses holistic rather than object-centric state representations.
Reinforcement Learning with Object-Centric Representations
@@ 6 Conclusion
We present SOLD, an object-centric model-based RL algorithm that learns directly from pixel inputs. By employing structured latent representations through slot-based dynamics models, our method offers a compelling alternative to traditional, holistic approaches. While object-centric representations have been valued for their role in forward prediction (Villar-Corrales et al., 2023), we demonstrate their synergistic benefits in accelerating the learning of behavior models. SOLD achieved strong performance across visual robotics environments, significantly outperforming the state-of-the-art DreamerV3, particularly in tasks requiring relational reasoning. Additionally, the learned behavior models exhibit interpretable attention patterns, explicitly focusing on task-relevant parts of the visual scene.
Limitations & Future Work
One limitation of our world model is that it generates predictions in a deterministic manner. This can be a drawback in environments that are inherently stochastic or highly unpredictable. We believe this is a key reason why SOLD outperforms DreamerV3 on complex robotic manipulation tasks but struggles to match its performance on simpler tasks like Cartpole-Balance, where minor variations in action sequences can lead to vastly different outcomes over long horizons. Addressing this limitation by incorporating stochasticity into the prediction model presents a promising direction for future work. A second limitation arises from the object-centric encoder-decoder model we use. While SAVi performs well on the tasks we evaluated, scaling it to complex real-world data remains a significant challenge. However, the core ideas of our method are independent of the specific object-centric encoder-decoder model, and future work can easily integrate more advanced models that address these scalability concerns.
@@ Acknowledgment
Reproducibility Statement
With the goal of reproducible research, we conducted three runs per method and task, each with different random seeds to account for seed-dependent variability. Additionally, we thoroughly list in Appendix B the model architectures and hyper-parameters used in our experiments. Upon acceptance of the paper, we will open-source our code, environments and pretrained models.
Appendix A Notation
@@
Appendix B Implementation Details
B.1 Slot Attention for Video
B.2 Object-centric dynamics model

### v2

@@ 4 Results
In this section, we present the empirical evaluation of SOLD on our suite of visual continuous control tasks. We first describe the comparative baselines and the environments used in our experiments. Using this setup, we aim to answer the following questions: (a) Does SOLD accurately model object-centric dynamics in action-conditional settings, preserving the decomposition of visual scenes? (b) Does the structured latent space allow SOLD to outperform SoTA baselines on tasks that require relation reasoning capabilities? (c) Can SOLD generalize to visual environments that are not object-centric by design?
Table 1: Final success rates. Success rates ($\pm$ standard deviation) of SOLD and baseline methods for the specific (top) and distinct (bottom) task variants.
(a) Specific tasks requiring mainly robotic control.
Task DreamerV3 TD-MPC2 w/o OCE SOLD
Reach $87.4$$\pm 1$ $97.6$$\pm 0$ $83.2$$\pm 2$ $\bm{97.9}$$\pm 0$
Reach-Rel. $45.6$$\pm 6$ $79.1$$\pm 1$ $39.2$$\pm 3$ $\bm{91.1}$$\pm 2$
Push $\bm{97.1}$$\pm 1$ $72.7$$\pm 3$ $75.2$$\pm 2$ $82.8$$\pm 2$
Pick $\bm{96.7}$$\pm 1$ $87.6$$\pm 2$ $22.9$$\pm 11$ $85.8$$\pm 7$
Average $81.7$$\pm 21$ $84.2$$\pm 9$ $55.1$$\pm 25$ $\bm{89.4}$$\pm 6$
(b) Distinct tasks requiring challenging relational reasoning.
Task DreamerV3 TD-MPC2 w/o OCE SOLD
Reach $14.6$$\pm 6$ $31.4$$\pm 3$ $11.3$$\pm 1$ $\bm{91.8}$$\pm 1$
Reach-Gr. $13.9$$\pm 2$ $15.7$$\pm 2$ $5.1$$\pm 1$ $\bm{69.6}$$\pm 2$
Push $70.0$$\pm 5$ $12.2$$\pm 5$ $10.5$$\pm 1$ $\bm{80.6}$$\pm 5$
Pick $33.9$$\pm 36$ $9.8$$\pm 1$ $0.7$$\pm 0$ $\bm{56.4}$$\pm 25$
Average $33.1$$\pm 23$ $17.3$$\pm 8$ $6.9$$\pm 4$ $\bm{74.6}$$\pm 13$
@@
The chosen baselines in our evaluation serve two primary purposes: assessing the impact of the object-centric paradigm in our method and benchmarking it against state-of-the-art approaches. To evaluate the effect of object-centric representations, we compare our method to a baseline that replaces the object-centric encoder-decoder modules with a standard convolutional architecture (w/o OCE). To benchmark against the best available methods, we include DreamerV3 (Hafner et al., 2023) and TD-MPC2 (Hansen et al., 2024). Both are widely recognized for their strong performance across a wide range of tasks. Additional details for the baselines can be found in Appendix E.
Figure 4: Achieved returns over the training duration across the eight benchmark environments. The dashed vertical line represents the offset for our method to account for the samples used during pre-training.
@@
We introduce a suite of eight object-centric robotic control environments designed to test both relational reasoning and manipulation capabilities. These environments feature two types of problems: Reach tasks, where the agent must identify a target and move the end-effector to its location, and manipulation tasks (Push and Pick), where the agent identifies a target block and moves it to a designated goal. To test varying levels of relational reasoning difficulty, we design the following configurations:
@@
For the Reach task, we include two advanced variants:
  • Specific-Relative    The goal is to reach the reddest object, determined by the perceptual CIEDE2000 (Sharma et al., 2005) distance.
  • Distinct-Groups    The environment contains 5 targets, and the goal is to reach the one that appears only once.
On these two additional reach tasks, we reuse the SAVi models that were pre-trained for Reach-Specific and Reach-Distinct, respectively without modification. Further details about these environments are provided in Appendix F.
@@ 4.1 Object-centric Dynamics Learning
The object-centric representations learned by SAVi can be seen in the context frame in Figure 3. The slots effectively decompose the visual scene, with most slots representing distinct objects and three slots capturing different parts of the robot. This part-whole segmentation highlights the ability of slots to meaningfully identify and represent separate parts of a larger object, such as the gripper jaws of the robot. Notably, the sharp mask predictions show that each slot isolates information about the specific object it represents (see also Section G). This property is crucial for object-centric behavior learning, as it enables subsequent components to reason about task-relevant objects while ignoring irrelevant information. Further, the open-loop prediction of 50 future frames starting from a single seed frame, shown in Figure 3, demonstrate the model’s ability to generate physically accurate predictions over long horizons. The movements of the robot and blocks are predicted with high accuracy, demonstrating the model’s ability to precisely capture physical interactions between objects. Moreover, the model effectively handles occlusions, as evidenced by the continued precise prediction of the spherical red target. Importantly, the autoregressive dynamics model maintains a precise and meaningful decomposition of the scene in its predictions, even far into the future. We encourage readers to view the videos of object-centric open-loop predictions on our project page for a qualitative assessment of these results.
@@ 4.2 Behavior Learning
To assess SOLD’s performance across our robotic control tasks, we train each method with three different random seeds per environment. The final success rates achieved by each method are presented in Table 1. SOLD consistently outperforms the non-object-centric baseline, often by a significant margin, underscoring the effectiveness of object-centric representations for the considered tasks. When compared to state-of-the-art MBRL methods, SOLD demonstrates competitive or superior performance. Notably, while SOLD narrowly surpasses DreamerV3 and TD-MPC2 across the Specific tasks, it significantly outperforms them on the more challenging Distinct variants, which require complex relational reasoning between objects. On these tasks, SOLD achieves more than double the performance of the second-best method.
Beyond that, when examining the performance over the course of training, as shown in Figure 4, we observe additional advantages in terms of sample efficiency. We find that SOLD consistently outperforms the highly sample-efficient DreamerV3 and TD-MPC2 baselines on all but the easiest Reach-Specific task, even after accounting for the samples used during pre-training. While the non-object-centric baseline demonstrates some success on the Specific tasks, it struggles with the relational reasoning required to solve the Distinct variants. In contrast, SOLD excels in tasks that demand reasoning about relationships between objects, as evidenced by the substantial performance gap observed on the Distinct tasks.
These results support our hypothesis that a structured latent representation within the world model significantly benefits tasks requiring object reasoning. This is especially valuable in robotics, where understanding object interactions is essential for solving complex control problems.
@@
To demonstrate that SOLD’s improved relational reasoning capabilities are accompanied by an interpretable focus on task-relevant objects, we visualize an excerpt of the slot history in Figure 5. To illustrate the actor’s attention pattern in the current (rightmost) time step, we multiply the attention scores by the masks of the respective objects and show them overlaid with the RGB reconstructions and as an individual colormap in the second and third rows, respectively. This visualization shows the Push-Specific task, where we find that the model automatically identifies task-relevant objects, disregarding slots that represent distractor objects across all time steps while focusing primarily on the robot and green cube. Although the recency bias induced by ALiBi is evident, we find that the model learns to overcome it when necessary, attending to the red sphere (indicating the goal position) after it has been occluded for 15 time steps, the last time it was visible. We see that the model effectively prioritizes task-relevant information, even when reasoning over long time sequences is required.
@@
A common limitation of prior work is the reliance on object-centric models pretrained on sequences with random behaviors and kept frozen during training, restricting their applicability to tasks where random and successful policies encounter similar state distributions. In the Pick tasks, this assumption is explicitly violated, as random behaviors rarely result in blocks being lifted off the table. Consequently, SAVi lacks prior exposure to configurations with blocks in the air, which are inevitable for successful policies. Figures 14 and 15 in the appendix illustrate the necessity of continually fine-tuning the object-centric encoder-decoder model: the fine-tuned model accurately reconstructs lifted blocks, whereas the frozen variant fails, causing blocks to dissolve during lifting.
4.3 Generalization to Non-Object-Centric Tasks
While object-centric methods are commonly evaluated on environments and datasets that naturally lend themselves to such decompositions, we aim to showcase the potential of our method to generalize beyond this setting. To this end, we train SOLD on the Button-Press and Hammer tasks from the Meta-World benchmark (Yu et al., 2019), both of which feature object with complex shapes and textures, challenging the model’s ability to handle more diverse and visually intricate inputs. Additionally, we test SOLD on the Cartpole-Balance and Finger-Spin environments from the DM-Control suite (Tassa et al., 2018), which represent significantly different domains not typically associated with object-centric learning. These environments are shown in Figure 6. SOLD achieves a $100\%$ success rate on both the Button-Press and Hammer tasks, highlighting its ability to adapt to visually diverse and challenging object interactions. On the Cartpole-Balance and Finger-Spin tasks, SOLD achieves returns of $497$ and $645$, respectively, demonstrating its capacity to generalize to tasks where object-centric reasoning is less pronounced. Details about environment decompositions and dynamics predictions for all four tasks can be found in Section G.2 of the Appendix.
[…]
Figure 6: Non-object centric environments from Meta-World (Yu et al., 2019) and DM-Control (Tassa et al., 2018).
@@ 5 Related Work
In recent years, the field of unsupervised object-centric representation learning from images and videos has gained significant attention (Yuan et al., 2023). Most existing methods follow an encoder-decoder framework with a structured bottleneck composed of $N$ latent vectors called slots, where each of these slots binds to a different object in the input image. Slot-based methods have been widely applied for images (Burgess et al., 2019; Locatello et al., 2020; Singh et al., 2021, 2023; Biza et al., 2023) and videos (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022; Bao et al., 2022). However, despite their impressive performance on synthetic datasets, they often fail to generalize to visually complex scenes. To overcome this limitation, recent methods propose using weak supervision (Elsayed et al., 2022; Bao et al., 2023), levering large pretrained encoders  (Seitzer et al., 2023; Aydemir et al., 2023; Kakogeorgiou et al., 2024), or using diffusion models as slot decoders (Jiang et al., 2023).
@@
Object-centric video prediction aims to understand the object dynamics in a video sequence with the goal of anticipating how these objects will move and interact with each other in future time steps. With this end, multiple methods propose to model and forecast the object dynamics using different architectures, including RNNs (Zoran et al., 2021; Nakano et al., 2023) transformers (Wu et al., 2023; Villar-Corrales et al., 2023; Song et al., 2023; Daniel & Tamar, 2024; Nguyen et al., 2024) or state-space models (Jiang et al., 2024), achieving an impressive prediction performance on synthetic video datasets and learning representations that can help solve downstream tasks that require reasoning about objects properties and relationships (Wu et al., 2023; Petri et al., 2024).
Model-based RL
Model-based RL approaches aim to improve sample efficiency by learning environment dynamics. PlaNet (Hafner et al., 2019) introduced a latent dynamics model for efficient planning, while the Dreamer family (Hafner et al., 2020, 2021, 2023) incorporated this into an actor-critic framework. DreamerV2 and DreamerV3 introduced further improvements like categorical latent states and robustness techniques. DreamerV3 has shown superior performance in visual control tasks compared to model-free approaches, but uses holistic rather than object-centric state representations. TD-MPC2 (Hansen et al., 2024), on the other hand, combines a reconstruction-free dynamics models with task-specific objectives and trajectory optimization, achieving strong performance in both state-based and visual control settings.
RL with Object-Centric Representations
@@ 6 Conclusion
We present SOLD, an object-centric model-based RL algorithm that learns directly from pixel inputs. By employing structured latent representations through slot-based dynamics models, our method offers a compelling alternative to traditional, holistic approaches. While object-centric representations have been valued for their role in forward prediction (Villar-Corrales et al., 2023), we demonstrate their synergistic benefits in accelerating the learning of behavior models. SOLD achieved strong performance across visual robotics environments, outperforming the state-of-the-art methods DreamerV3 and TD-MPC2, particularly in tasks requiring relational reasoning. Additionally, the learned behavior models exhibit interpretable attention patterns, explicitly focusing on task-relevant parts of the visual scene.
@@ Acknowledgment
Appendix A Limitations & Future Work
One limitation of our world model is that it generates predictions in a deterministic manner. This can be a drawback in environments that are inherently stochastic or highly unpredictable. Addressing this limitation by incorporating stochasticity into the prediction model presents a promising direction for future work. A second limitation arises from the object-centric encoder-decoder model we use. While SAVi performs well on the tasks we evaluated, scaling it to complex real-world data remains a significant challenge. However, the core ideas of our method are independent of the specific object-centric encoder-decoder model, and future work can easily integrate more advanced models that address these scalability concerns.
Appendix B Notation
@@
Appendix C Extended Related Work
Model-based Reinforcement Learning
Model-based methods hold the potential to significantly improve the sample efficiency of RL algorithms, and recent years have seen several key contributions advancing this area.
Pioneering work by Ha & Schmidhuber (2018) introduced the concept of a recurrent generative model, termed a world model, which captures the dynamics of visual RL environments. By encoding high-dimensional observations into a compact latent representation, this model enables RL agents to train policies entirely within imagined rollouts.
The Planning Network (PlaNet) (Hafner et al., 2019) introduced a recurrent state-space model (RSSM) that predicts future states directly in a compact latent space, avoiding the costly step of decoding full observations. This architecture enables efficient planning of action sequences but is limited by short planning horizons. Building on this, Dreamer (Hafner et al., 2020) integrates planning and learning by training agents within a learned world model, overcoming PlaNet’s shortsighted horizon. Subsequent versions, DreamerV2 (Hafner et al., 2021) and DreamerV3 (Hafner et al., 2023), improved robustness and generalization through enhanced representation learning and optimization techniques, achieving state-of-the-art results across diverse RL benchmarks.
Temporal Difference Learning for Model Predictive Control (TD-MPC) (Hansen et al., 2022) introduced a task-oriented latent dynamics model to optimize trajectories directly within the latent space of a world model. Unlike earlier approaches, TD-MPC avoids reconstructing full observations, instead focusing the world model on reward-predictive elements through a loss applied to reward and value predictions. TD-MPC2 (Hansen et al., 2024) builds on this by introducing scalability improvements, enabling superior performance with larger model sizes and demonstrating a single agent’s ability to generalize across multiple tasks and action spaces.
Inspired by the success of Transformers on sequence modeling tasks, Micheli et al. (2023) proposed IRIS, a method combining a discrete autoencoder with an autoregressive Transformer to model environment dynamics. The autoencoder tokenizes images into a discrete set of representations, while the Transformer learns temporal dynamics across these tokens. IRIS demonstrated visually and temporally accurate predictions of game dynamics in Atari environments. While IRIS shares similarities with our approach – encoding an image into a set-based representation and predicting it forward using a Transformer – it lacks the object-centric interpretability afforded by our model.
Recently, Alonso et al. (2024) proposed diffusion as a promising alternative to discretization of the latent space of the world model. DIAMOND uses a diffusion-based conditional generative model, $p(\bm{x}_{t+1}\mid\bm{x}_{\leq t},\bm{a}_{\leq t})$, to produce visually precise next-frame predictions. The authors demonstrate the potential of their world model to simulate complex 3D environments by learning a realistic game-engine from static Counter-Strike: Global Offensive gameplay. Valevski et al. (2024) also propose to use diffusion to create high-quality visual predictions. Specifically, they demonstrate the potential of diffusion models to serve as real-time game engines.
Appendix D Implementation Details